You decide to invest money in a T-bill. A T-bill is a bond that is sold at a discount face value. For example, suppose you buy a 13 week T-bill with a face value of $10,000 for $9,800. What annual simple interest rate have you earned?
Answer #1

Solution :

13-week T-bill will get 10000-9800 = 200

So an year has 52 weeks

interest for 13 week period = 200/9800 = 0.02040816 for 4 times an year

nominal interest rate per year = 0.02040816 * 4 = 0.0816326

0.0816326 * 100 =  8.16%

Annual interest rate =  8.16%


If you liked the answer, please give an upvote. This will be quite encouraging for me, thank-you!